1st November 2006 Hoyle Committee Room 10:00 - 12:37 Present: DWE, MJI, JRL, PSB, STH, EGS, WJS, MR Apologies: RGM Guest Appearance: NAW Agenda ------ 1. Actions from last meeting 2. Comments on WFAU minutes 3. Recent meetings - ADASS 2006 Tucson - UKIDSS meeting in Nottingham - VDFS UK review 4. Data archives update - AAT, ING, WFCAM, UKIRT 5. Optical processing update- IPHAS release via AG tools 6. WFCAM news - 06B news - photometric recalibration 7. VISTA news - VIRCAM lab data analysis - CASU & commissioning tasks - VISTA survey management plan & public surveys 8. AOB Minutes ------- 1. Actions from last meeting JRL has nearly implemented the fix to the problem with the UKIRT archive (missing header information problem). Only a few modifications to scripts need doing - ongoing <<<< STH has started combining the data of the supernova remnant image from Albert Zijlstra to make a nice colour image. He has now passed on the task to Amanda Smith. He has also emailed ING to see if we can obtain their poster directly - ongoing <<<< STH has not investigated the effects of stacking on the 2MASS/WFCAM non-linearity comparison - ongoing <<<< JRL has come up with a solution regarding the PostgreSQL problems with the AAT archive. However, the communication rate is slow and holding up the completion of this task - ongoing <<<< On a related topic, it was pointed out that it has been quite a while since any AAT data has been received by CASU. The backlog of data not received will stretch to 2 years soon. MJI have not emailed the new WFCAM colour equations to DWE and JRL (see STH next action) - ongoing <<<< STH has discussed the new WFCAM colour equations within the Calibration working group. A calibration scheme using the JHK colour diagram to calibrate Y and Z is favoured and will be agreed soon - ongoing <<<< RGM has not finished working on the classification aspect of cross-talk weeding - ongoing <<<< JRL has not added a detection filter FWHM choice as a command line option of the pipeline version of imcore. There was some discussion regarding this and that exact task has been clarified - ongoing <<<< ALL sent in quarterly reports to STH. In the end, these were not needed for the UK Review. DWE have not updated the Plone technical pages on astrometry. At the moment, MJI the description in the pages agrees with the released data. When the astrometry gets updated, then the pages will be updated - ongoing <<<< MJI raised the issue of UKIDSS data release policy at the Nottingham meeting. Apparently the JAC web page is not meant to say that the data will be released 18 months after observation. It should say 18 months after the processed data has been released to the European community. SJW will contact AA regarding this. PSB requested the latest official VISTA schedule and distributed it ALL suggested suitable venues for the UK Review meal. An excellent meal at Fitzbillies was had. 2. Comments on WFAU minutes The deadline of 25 October was noted for gripes from the Survey Heads regarding 05A reductions. MJI said that only one had come to his attention, concerning problems with the crosstalk correction in areas of bright nebulosity. A logic bug has been detected and a fix is being worked on. It was pointed out that data with the crosstalk correction has been available for over a year. MJI has written a script for stripping out parts of the FITS headers and putting them in a file. This has been sent to WFAU. 3. Recent meetings JRL and PSB reported on items of interest from the ADASS 2006 meeting in Tucson. FITS: the proposal to have header records of variable length was not supported. Neither was that to have keywords longer than 8 characters. WCS: there was some progress with the definition of distortions that aren't described by the usual functional forms. This will be done via tables. Intrapixel losses in photometry: there was a good talk on this by Kenneth Mighell. MJI asked if anyone was interested in investigating this. MJI said that he would email Derek Ives to see if there was any WFCAM information on this. <<<< (There was also mention of the QFITS/CFITSIO using groups in ESO merging and of a talk on VISTA quality control. Not sure if these need mentioning) MJI attended an amazingly boring :) UKIDSS meeting in Nottingham on 13 October. Most of the discussion was about the new proposals and very little on the VDFS pipeline. The old surveys will be resubmitted along with one new one (UHS - UKIRT Hemisphere Survey). There were many queries about where the extra observing time would be coming from. This has implications for CASU in that there will be more processing to be done. The UK VDFS review was held on 24/25 October at Cambridge. This was a hugely enjoyable and worthwhile meeting. :) :) Colin Vincent from PPARC attended some of the meeting which was good, since this reminds the funding agency where the money is going. At some point we will receive a written reply/report. The immediate verbal response was positive. A suggestion from the review panel was that the software be made available to the public. The CASU view is that this will invariably involve supporting the code and a consequent use of CASU resources. 4. Data archives update AAT: No further news ING: This is now stable. EGS was asked to make a backup of the archive. <<<< MJI said that new LTO tapes have been bought for the purpose. JRL raised the issue of migrating the Sybase system off the Suns and onto apm20. This was because we want to retire cass36 and cass38 by the end of the year. There was some discussion regarding the VizieR UK mirror that we host and how it would be affected by this. WFCAM: Is up-to-date. WFCAM is now back on the telescope and that we expect data in the near future. UKIRT: Is up-to-date. 5. Optical processing update EGS said that the IPHAS release (R, I and H_alpha) will be made available to the community via AstroGrid tools. He pointed out that only one piece of software needs to be downloaded in order to use this. Initially 120 square degrees were released in order to test the system and receive comments. Since there were no adverse comments, the whole survey is now being ingested. However, there are a few calibration issues that need sorting out. It was noted that more IPHAS data is being taken. EGS added that CMC14 was also being released via AstroGrid. The various rumours about the VST were discussed. 6. WFCAM news 06B news: MJI reported on the latest from WFCAM which was now back on the telescope. Tests are underway on the new baffling tube which should cut out the Moon glint. New acquisition hardware has been installed which should reduce the number of detector dropouts that are experienced. A new observing strategy will also be tested out. New tapes are expected in about a week. It was noted that the photometric recalibration was needed for the DR2 release and was thus needed as soon as possible. PSF fitting: DWE said that the remaining tasks with PSF fitting were to implement a variable cutout and optimize the background estimation. He was asked to email Phil Lucas in order to progress comparisons with other <<<< software packages. GPS and M17 fields would be suitable for this. It was noted that the data transfers have stopped. MJI reported a conversation with NCH about automating the transfers which NCH agreed was possible. PSB said that over the last quarter, the IoA was the department with the largest amount of network traffic. MJI drew our attention to the advert for a UKIRT survey manager. 7. VISTA news WJS reported on his week long stay at Paranal. The general impression was good. The dome was almost light tight and the mechanics and electronics of the telescope seem OK. There were a few niggles about the mirror supports, but were OK. However, if there is a software crash, the mirror is dropped! (by 1mm). The mirror polishing remains on the critical path. The acceptance in Moscow is scheduled for December, but January is more likely. The camera is in good shape and will ship from RAL by the end of the year. The main thing holding this up is the requirement that the documentation gets shipped at the same time. Commissioning is expected in Autumn 2007. There was further discussion regarding the CASU involvement in the commissioning. There was some discussion regarding the purchasing of computer hardware for Paranal. It was felt that this should be done soon. MJI said that he would find out from JPE how his meeting went with Michele Peron. <<<< MJI said that the PSP meeting is being held around now to decide on the public surveys. After this a new management plan for the reductions will have to be drawn up. This will be around December and the timescale will be quite tight. JRL reported on the VIRCAM AIT dark frames. There are many hot pixels, which all have a cross shape. Dark subtraction will remove the transepts leaving single hot pixels. These cannot be flat fielded out and some chips have about 1% of these remaining hot pixels. Some of these hot pixels are in clusters. In order to carry out the dark correction you need a dark with a matching exposure time otherwise the transepts are not removed. Curtaining is present in VISTA data, but at a lower level than for WFCAM. This goes across channels and all 4 detectors and is thus caused by a single controller. This can thus be measured by taking a median along the row across all channels and the 4 detectors. This pattern is different between the 4 controllers. This will not be removed in the summit pipeline. MJI said that he would contact Mark Casali about the controllers since they are ESO ones. <<<< There is a dead patch on Detector 1 which does subtract out, however, nothing can be done with this section ("It's dead Jim"). The dark current is quite high at 0.5 ADU/s/pixel. The read out noise is about 25 e- and the gain around 4-5 e-/ADU. A worry is that on Detector 14 a complete channel keeps dropping in and out. JRL will release a report on this soon. <<<< 8. AOB MJI said that the registration deadline for the ESO Calibration Workshop to be held at Garching is 6 November. PSB, MJI, MR and JRL will attend. The central heating in the APM building is finally working properly. The warped windows have been bodged shut for the winter. Presumably they will be bodged open again for the summer. FvL wants to know about who wants to work for Gaia fairly soon (within a week?). This is so that the amount of new recruitment can be worked out. All were asked to think about this and then go and chat with FvL and MJI. <<<< NAW raised the issue of buying hardware to run the service distributing IPHAS data. There was some discussion as to who pays for this and how powerful a machine was required. It was decided that a machine would be bought as a spare for running the CASU web services which could be used for the IPHAS service in the meantime. ???? Continuing Actions ------------------ JRL finish the fix to problem with UKIRT archive (missing header information) STH generate poster of supernova remnant image received from Albert Zijlstra STH investigate effects of stacking on 2MASS/WFCAM non-linearity comparison JRL repair AAT archive mirroring (PostgreSQL problems) MJI email new WFCAM colour equations to DWE and JRL STH STH agree new WFCAM colour equations with Calibration working group RGM report back on classification aspect of cross-talk weeding JRL add detection filter FWHM choice as command line option of pipeline version of imcore. DWE update Plone technical pages on astrometry when astrometry gets updated MJI New Actions ----------- MJI email Derek Ives to find out about WFCAM intrapixel losses EGS backup ING archive DWE contact Phil Lucas regarding PSF fitting comparisons MJI find out from JPE about meeting with Michele Peron MJI contact Mark Casali about curtaining in ESO controllers JRL release report on VIRCAM AIT dark frames ALL think about Gaia involvement and discuss with MJI and FvL ??? purchase new server as spare for CASU web services and for running IPHAS service